OpenStudy (ghosting):

I feel like this question is probably actually really simple, and I am just not understanding it... But I need a little help, please! It says: This expression evaluates to an integer. Write the integer in the space below, like this: -14 Then it has these numbers... 3 5/3 (yeah a fraction right here) ------------ (then this line) 3 2/3 (and another fraction here) There is also a line to the right side of the problem. Unsure what this line represents... Could someone explain some of this to me?

OpenStudy (johnweldon1993):

So it looks like \(\Large \frac{3\frac{5}{3}}{3\frac{2}{3}}\) ?

OpenStudy (johnweldon1993):

Nope cant be that, that's not an integer...

OpenStudy (ghosting):

That is how it looks though!

OpenStudy (ghosting):

But with another line vertically next to the problem, to the right of it.

OpenStudy (ghosting):

like this | but longer

OpenStudy (johnweldon1993):

Hmm, any way you can take a screenshot of it and attach it here?

OpenStudy (ghosting):

Yeah gimmie a second!

OpenStudy (jtug6):

Ah I see what is being asked. It's the fractions separated and they want it in its simplified integer form. Just simplifying both fractions

OpenStudy (johnweldon1993):

Take it away jtug!

OpenStudy (johnweldon1993):

Oh now I get it....

OpenStudy (ghosting):

Okay, that makes sense kinda... But, I am a bit confused on how to simplify these together for one integer... How should I do this??

OpenStudy (jtug6):

Take the denominator multiply it by the whole number to the left then add it to the numerator

OpenStudy (johnweldon1993):

No, dont fall into that trap, thats what I thought too...but no these are not mixed numbers, we are dealing with exponents I had \(\Large \frac{3\frac{5}{3}}{3\frac{2}{3}}\) when it should be \(\Large \frac{3^\frac{5}{3}}{3^\frac{2}{3}}\)

OpenStudy (johnweldon1993):

From this, it is important to remember the laws of exponents \[\large \frac{x^a}{x^b} = x^{a - b}\]

OpenStudy (jtug6):

Oh! I didn't look @ the image. Okay so it's an exponent? That makes sense then

OpenStudy (jtug6):

My bad thought it was two separate mixed numbers XD

OpenStudy (johnweldon1993):

You with us @ghosting ?

OpenStudy (ghosting):

Nope, I am long gone!!! @__@ What do I do now.... confuzzled, sorry.

OpenStudy (ghosting):

Okay so... those fractions are actually squared by the whole numbers... I get that part.. .__. I think. Right?

OpenStudy (johnweldon1993):

Nope dont let it overwhelm you! Remember an exponent is simply something raised to a power \(\large 2^4, 6^{10}, 4^{22}\) etc... So let me take a tangent quick here... If I gave you \(\Large \frac{2^4}{2^2}\) and told you the law to follow is \(\Large \frac{x^a}{x^b} = x^{a - b}\) What would you get?

OpenStudy (ghosting):

4?

OpenStudy (johnweldon1993):

That is the correct answer yes...and you got it by doing \[\large \frac{2^4}{2^2} = 2^{4 - 2} = 2^2 = 4\] Right?

OpenStudy (ghosting):

Correct!

OpenStudy (johnweldon1993):

Good! So back to your problem \[\large \frac{3^\frac{5}{3}}{3^\frac{2}{3}}\] Treat it the SAME way..doesn't matter that we are working with fractions this time

OpenStudy (ghosting):

I have no idea what to do with the fractions, though. Would I multiply 5/3 3 times? or..

OpenStudy (johnweldon1993):

Nope...just like the last problem...you left the big 2 alone right? And only subtracted the exponents (4 - 2) Here you do the same...leave the big 3 alone...and subtract the exponents \[\Large \frac{3^\frac{5}{3}}{3^\frac{2}{3}} = 3^{\frac{5}{3} - \frac{2}{3}}\] So just ask yourself...what is \(\large \frac{5}{3} - \frac{2}{3}\)

OpenStudy (ghosting):

We get 1. :D so it's 3^1? so just 3, right?

OpenStudy (johnweldon1993):

Exactly! :)

OpenStudy (ghosting):

COOL, Wow! Thank you so much for the help!

OpenStudy (johnweldon1993):

No problem!